When 1 mole of this fluorocarbon compound was completely combusted in a bomb

calorimeter, it liberates 3870 kJ of heat which causes the temperature to increase from

25 °C to 26.5 °C. Compute the heat capacity of the bomb calorimeter?


1
Expert's answer
2021-12-16T11:09:04-0500

qcal = Ccal (T2-T1)

qcal = -qrxn

Ccal = -qrxn / (T2-T1) = -(-3870 kJ) / (26.5 - 25) = 2580 kJ/0C